(D² +40D'² +3D' )Z=ycos x

1
Expert's answer
2022-01-31T16:31:50-0500

The actual question is;

"\displaystyle
(D^2+4DD^\prime+3(D^\prime)^2)z=y\cos x, \text{where }D=\frac{\partial}{\partial x}, \text{and}\ D^\prime =\frac{\partial}{\partial y}"


To find the complementary function (C.F), let "\displaystyle
D=m, \text{and }D^\prime=1," then the auxiliary equation of the given PDE in terms of variable m is;

"\displaystyle
m^2+4m+3=0"

"\displaystyle
\Rightarrow (m+1)(m+3)=0\Rightarrow m=-1,-3."

Thus the complementary function of the given PDE is;

"\displaystyle
\color{green}{C.F=z_c=f_1(y-x)+f_2(y-3x)}"


Next, to find the particular integral (P.I) of the given PDE, we proceed as follows;

"\displaystyle
P.I=\frac{1}{(D^2+4DD^\prime+3(D^\prime)^2)}y\cos x=\frac{1}{(D+D^\prime)(D+3D^\prime)}y\cos x"

"\displaystyle
\ =\frac{1}{(D+D^\prime)}\int (c+3x)\cos x\ dx, \text{where }y=c+3x\\"

"\displaystyle
\ =\frac{1}{(D+D^\prime)}[(c+3x)\sin x+3\cos x], \text{ by integration by parts}"

"\displaystyle
\ =\frac{1}{(D+D^\prime)}[y\sin x+3\cos x], \text{ since }y=c+3x"

"\displaystyle
\ =\int[(c+x)\sin x+3\cos x]\ dx, \text{where we put }y=c+x"

"\displaystyle
\ =\int[(c+x)\sin x\ dx]+3\int \cos x\ dx"

"\displaystyle
\ =[(c+x)(-\cos x)+\sin x]+3\sin x\ ,\text{integration by parts}"

"\displaystyle
\ =-y\cos x+4\sin x,\ \text{where we put back }y=c+x"

Thus the particular integral of the given PDE is;

"\displaystyle
\color{blue}{P.I=z_p=-y\cos x+4\sin x}"

Since complete solution (z) is "\displaystyle
\color{red}{z=z_c+z_p}", thus we have the complete solution as;

"\displaystyle
\color{red}{z=f_1(y-x)+f_2(y-3x)-y\cos x+4\sin x}"